OpenStudy (anonymous):

Find the 7th term of the expansion of (3c + 2d)^9

OpenStudy (campbell_st):

so using the binomial expansion its \[^9C_{6} (3c)^{9 - 6}(2a)^6\] just evaluate

OpenStudy (anonymous):

How do I evaluate that?

OpenStudy (campbell_st):

well \[^9C_{6}\] is a combination which can be evaluated using a calculator so you have \[^9C_{6} \times (3)^3 \times (2)^6 \times(c)^3 \times (d)^6\] hope this makes it easier

OpenStudy (anonymous):

Which one of these would it be?? a.760c3d6 c.145,152c3d6 b.760c4d5 d.145,152c4d5

OpenStudy (anonymous):

I think its B..

OpenStudy (campbell_st):

lol... well its A or B... what is 3^3 x 2^6 = that should help eliminate one choice...

OpenStudy (anonymous):

108